Solve for y
x+5y=15
Hint: Isolate y on one side of the equal sign.

Answer by stanbon(48558)

Solve for y
x+5y=15
5y=15-x
y=(15-x)/5
y=3-(1/5)x
=========
Cheers,
Stan H.